OpenStudy (anonymous):

Make "v" the subject of the formula: m = m / (sqrt 1 - (v/c)^2 ) Lol kinda eh confusing...just hmu n I'll clarify anything.

OpenStudy (anonymous):

from looking at it v should equal 0 but lets go thru it divide each side by m: 1=1/(sqrt(1-(v/c)^2)) multiply each side by the denominator on the right: sqrt(1-(v/c)^2)=1 square both sides 1-(v/c)^2=1 subtract a 1 from each side -(v/c)^2=0 multiply each side by (-1) (v/c)^2=0 square root both sides v/c=0 this can only occur if v=0 and \[c \neq0\]

OpenStudy (anonymous):

Makes sense Jah. Thank you. :)